# Nature of spinons in the 1D spin chains

### v2

@@ Abstract
Figure 1: Cartoon views of collective excitations in spin-$\frac{1}{2}$ Heisenberg chains. Panel (a): canonical picture arising after flipping one spin in the Heisenberg FM ground state (left subpanel) and triggering a propagating magnon (middle subpanel) which is understood as a free mobile particle in the vacuum (right subpanel). Panel (b): canonical picture arising after flipping one spin in the Heisenberg AF ground state (left subpanel) and triggering two propagating spinons (middle subpanel) which are usually approximately understood as two mobile domain walls in a NĂŠel AF (right subpanel). Panel (c): picture introduced in this paper, arising after inserting an additional spin in the Heisenberg AF ground state (left subpanel) and triggering a propagating spinon (middle subpanel) which can be understood as a mobile spin flowing in the valence-bond solid (right subpanel). Flipped or inserted spins are shown in red; entanglement between two or more sites is shown by a blue-coloured hatched area.
@@
An important feature of the construction, further discussed below, is that $|\Psi(m)\rangle$ is not normalized. Denoting by $|\bar{\Psi}(m)\rangle$ its normalized counterpart, we find the first two moments of the local spin operators to be ⟨$\langle\bar{\Psi}(m)|\sigma_{j}^{z}|\bar{\Psi}(m)\rangle=\frac{\delta_{jm}}{2}$, ⟨$\langle\bar{\Psi}(m)|(\sigma_{m}^{z})^{2}|\bar{\Psi}(m)\rangle=\frac{1}{4}$, and for $j\neq m$: ⟨$\langle\bar{\Psi}(m)|(\sigma_{j}^{z})^{2}|\bar{\Psi}(m)\rangle=\langle\Psi_{0}|(\sigma_{1}^{z})^{2}|\Psi_{0}\rangle$, where in the last equality we used the translational invariance of the ground state. This result can be visualized as a spin up frozen at the $m^{th}$ site of the $L+1$-site chain, in the sea of fluctuating spins on all other sites. These fluctuations are identical to the ground state fluctuations of the $L$-site chain, by construction. This picture is in accordance with the concept of elementary excitations which are created from the ground state, and inherit rather than rebuild the correlations present in it.
To define a dispersion relation we introduce a Fourier transform with momentum $q=\frac{2\pi n}{L+1}$, −$-L/2\leqslant n\leqslant L/2$, $n\in\mathbb{Z}$:
@@
In our procedure, the correlations present in the spinon state are inherited from the ground state. This leads to the natural question as to which correlations, of all those present in the exact ground state, are most relevant to the spinon. To answer this question we consider two approximations to the ground state: (i) the lowest energy classical configuration, i.e. the NĂŠel state, and (ii) the simplest valence-bond solid.
@@
The corresponding $\varepsilon(q)$, $N(q)$ are shown in Fig. 2(c1-c2). The dispersion relation is qualitatively similar to the spinon dispersion (black solid line). Quantitatively, the value at $q=0$ is underestimated and the value at $q=\frac{\pi}{2}$ is negative, reflecting the fact that the valence-bond state is an approximation. Indeed, its average energy per site is −$-0.3750$, significantly larger than the exact ground state energy per site of −$-0.4438$ Auerbach (1994).
@@
Dispersion linear in momentum: The other notable feature of the spinon dispersion is its linear dependence on $q$ in the low-energy limit. It can be understood as a result of (i) the vanishing compact support for $|q|>\pi/2$ that is due to ground state entanglement (see above), and (ii) the ∝$\propto\cos q$ dispersion relation arising from nearest neighbor hopping of the spinon, discussed next.
@@
Outlook: The postulated procedure is universal and can be applied to any 1D spin model, cf. Sec. III of SM for a short overview. In particular, it can be used to study excitations of the 1D FM chain, see Fig. 2(a-b) of SM : in this case a single magnon is excited in the procedure. This further shows that the spinon is not exotic because it can be excited in quite a special manner but it is the entanglement nature of the antiferromagnetic ground state that is crucial here. We suggest that in the future, this procedure can be used to explore the nature of the elementary excitations in frustrated spin systems, such as the recently discussed problem of marginally irrelevant interactions between spinons in $J_{1}-J_{2}$ spin chains Keselman et al. (2020). We propose to generalise the procedure to higher dimensions by focusing on the Green’s function of an immobile hole added to a spin system, for this corresponds to removing a spin from a lattice.
@@
Acknowledgements.
Authors acknowledge the support of National Science Centre in Poland under the projects 2016/22/E/ST3/00560 (K.W.), 2018/31/B/ST3/03758 (T.K.), and 2018/31/D/ST3/03588 (M.P.). K.W. acknowledges support from the Excellence Initiative of the University of Warsaw (the ‘mikrogrant’ programme). M.B. acknowledges support from the Stewart Blusson Quantum Matter Institute and from NSERC. We thank Federico Becca, Akira Furusaki and Karlo Penc for stimulating discussions and all the organisers of the KITP program “A New Spin on Quantum Mangets” for allowing us to present this work during the program. We thank the Referee for bringing Ref. Talstra and Strong (1997) to our attention, Karlo Penc for pointing out references  Sorella and Parola (1992); Penc et al. (1997); Penc and Serhan (1997), and Akira Furusaki for pointing out references  Matveev et al. (2007a, b). This research was supported in part by the National Science Foundation under Grant Nos. NSF PHY-1748958 and NSF PHY-2309135. This research was carried out with the support of the Interdisciplinary Centre for Mathematical and Computational Modelling at University of Warsaw (ICM UW) under grant no. G73-29.
@@ I Exact result for the $4$-site Heisenberg chain: relation between the norm $N(q)$ and entanglement
$|\Psi_{0}\rangle=\frac{|\!\uparrow\downarrow\rangle-|\!\downarrow\uparrow\rangle}{\sqrt{2}}\otimes\frac{|\!\uparrow\downarrow\rangle-|\!\downarrow\uparrow\rangle}{\sqrt{2}}+{\rm(translation)}.$ (11)
@@
$|\Psi(q)\rangle=\frac{1+e^{iq}-\frac{e^{3iq}}{2}}{\sqrt{3}}|A\rangle-\frac{1+e^{iq}+e^{2iq}}{2\sqrt{3}}|B\rangle$ (12)
with $|A\rangle\!\!\!=\!\!\!|\!\!\uparrow\uparrow\downarrow\uparrow\downarrow\rangle\!\!+\!\!({\rm translations})$, $|B\rangle\!\!\!=\!|\!\!\uparrow\uparrow\uparrow\downarrow\downarrow\rangle+({\rm translations})$. All the states in $|A\rangle$ and $|B\rangle$ are orthonormal to each other. The resulting norm:
$N(q)=1+\cos q-\frac{1}{6}\cos(2q)-\frac{1}{3}\cos(3q)$ (13)
@@
Having obtained an exact expression for $N(q)$, (13), we are now ready to investigate the origin of its two main striking properties: (i) a relative enhancement of the $4$-site Heisenberg chain norm, with a pronounced maximum for the small momenta $q\lesssim\pi/2$, w.r.t. the norm obtained once a spin is added to the Neel antiferromagnet, that is w.r.t. $N(q)=1+\cos(q)$ (see main text); (ii) a relative suppression of the $4$-site Heisenberg chain norm, with a pronounced minimum for the small momenta $q\gtrsim\pi/2$, w.r.t. the norm obtained once a spin is added to the Neel antiferromagnet, that is w.r.t. $N(q)=1+\cos(q)$ (see main text).
To this end, let us first note that a sum rule ∑$\sum_{0<q<\pi}N(q)=\pi$ means that once we e.g. understand the relative enhancement for some momenta, then the relative suppression for the other momenta follows. Therefore, we concentrate on investigating the former effect.
A quick inspection of Eq. (13) shows that the term responsible for obtaining such a relative enhancement in $N(q)$ for small momenta (and, conversely, also a relative suppression for large momenta) is the term ∝−$\propto-\cos(3q)$. A closer look at the procedure of adding an extra spin shows that, in order to obtain such a contribution to the norm, there must exist two product states in the ground state which contribute to $|\Psi_{0}\rangle$ with opposite sign and that they must be identical to each other after adding an extra spin in our procedure, modulo a phase factor $e^{3iq}$. This is fulfilled for instance by such two contributions to the exact ground state $|\Psi_{0}\rangle$ that, upon adding an extra spin site at particular location, lead to the same states which differ by $e^{3iq}$ in our procedure:
@@
−$-\frac{1}{2}|\uparrow\uparrow\downarrow\downarrow\rangle\quad$−$\rightarrow\quad-\frac{1}{2}e^{3iq}|\uparrow\uparrow\downarrow\tilde{\uparrow}\downarrow\rangle,$ (15)
@@
Identifying the type of product states in $|\Psi_{0}\rangle$ that lead to the relative enhancement of the norm $N(q)$ for small momenta, allows us to intuitively understand the origin of this phenomenon: namely, it follows from the valence bond nature of the ground state—the two ‘fluctuating singlets’ allow for the onset of ferromagnetic domains with opposite magnetisation and a (relative) negative phase contribution to the ground state, see Eq. (15). Note that the relative negative phase is just as crucial as the existence of ferromagnetic domains and that such a phase difference can only be realised once $|\Psi_{0}\rangle$ is not a product state. Altogether, this shows how specific properties of the norm $N(q)$ of the state constructed by adding one extra spin can be related to the entangled nature of the spin chain ground state $|\Psi_{0}\rangle$.
@@ II The norm $N(0)$ as an entanglement witness
$|\Psi(q=0)\rangle=\frac{1}{\sqrt{L+1}}\bar{C}_{L}|\Psi_{0}\rangle.$ (18)
@@
$N(0)=\frac{1}{L+1}\frac{\langle\Psi_{0}|\bar{C}_{L}^{\dagger}\bar{C}_{L}|\Psi_{0}\rangle}{\langle\Psi_{0}|\Psi_{0}\rangle},$ (19)
@@
$T_{L}|\Psi_{1}\rangle=T_{L}|\Psi_{2}\rangle,\qquad|\Psi_{i}\rangle\in[|\Psi_{0}\rangle].$ (20)
@@
$\mathcal{N}_{\Psi_{0}}=\frac{1}{L+1}\frac{\langle\Psi|T_{L}^{\dagger}\bar{C}_{L}^{\dagger}\bar{C}_{L}T_{L}|\Psi\rangle}{\langle\Psi|T_{L}^{\dagger}T_{L}|\Psi\rangle}.$ (21)
From the property (20) it follows that $\mathcal{N}$ is the same for any $|\Psi\rangle$ in the equivalance class $[|\Psi_{0}\rangle]$ of $|\Psi_{0}\rangle$. This implies that to compute $\mathcal{N}$ for a translationally invariant state we can identify one of its seeds and then use equation (21). This equation still explicitly symmetrizes over all translations but we will now rewrite in a form more suitable for further computations. In this process the above mentioned symmetry becomes less obvious.
@@
Using the properties of operators $C_{L}$ and $T_{L}$ we can prove a simple property of ⟨$\langle\Psi_{0}|\bar{C}_{L}^{\dagger}\bar{C}_{L}|\Psi_{0}\rangle$ that, due to the translational invariance, position of one of the insertions can be fixed,
⟨$\langle\Psi_{0}|\bar{C}_{L}^{\dagger}\bar{C}_{L}|\Psi_{0}\rangle$ =⟨$=\langle\Psi|T_{L}\bar{C}_{L}^{\dagger}\bar{C}_{L}T_{L}|\Psi\rangle$
=⟨$=\langle\Psi|C_{L}^{\dagger}T_{L+1}T_{L+1}c_{m}T_{L}|\Psi\rangle$
$=(L+1)\langle\Psi|\bar{C}_{L}^{\dagger}T_{L+1}c_{m}T_{L}|\Psi\rangle.$ (22)
@@
$\mathcal{N}=\frac{1}{L}\frac{\langle\Psi|T_{L}\bar{C}_{L}^{\dagger}c_{0}T_{L}|\Psi\rangle}{\langle\Psi|T_{L}|\Psi\rangle}.$ (23)
@@
$|\Psi\rangle=\bigotimes_{i=1}^{k}|\psi\rangle,$ (25)
@@
$\mathcal{N}_{\Psi_{0}}\geq N_{\psi_{0}}.$ (26)
Assuming (24) we have $T_{L}|\Psi\rangle=kT_{m}|\Psi\rangle$. We then find
@@
≥$\geq\frac{1}{m}\frac{\langle\Psi|T_{m}\bar{C}_{m}^{\dagger}c_{0}T_{m}|\Psi\rangle}{\langle\Psi|\Psi\rangle}=\frac{1}{m}\frac{\langle\psi|T_{m}\bar{C}_{m}^{\dagger}c_{0}T_{m}|\psi\rangle}{\langle\psi|\psi\rangle}=\mathcal{N}_{\psi_{0}},$ (27)
@@
This proves that $\mathcal{N}$ for a seed $|\Psi\rangle$ which is a product state can be bounded by $\mathcal{N}$ of the elementary cell $|\psi\rangle$ of the seed. This is important because if we now want to find further bounds on $\mathcal{N}$ we can focus on states which do not have any invariance under translations which simplifies the analysis. To this end, we observe that for state $|\Psi_{0}\rangle$ such that its seed $|\Psi\rangle$ does not have any translational symmetry the formula (21) can be reduced to
$\mathcal{N}_{\Psi_{0}}=\frac{1}{L}\frac{\langle\Psi|C_{L}^{\dagger}T_{L+1}C_{L}|\Psi\rangle}{\langle\Psi|\Psi\rangle},$ (28)
where we used $T_{L}^{2}=LT_{L}$ and due to lack of the translational symmetry ⟨$\langle\Psi|T_{L}|\Psi\rangle=\langle\Psi|\Psi\rangle$.
@@
⟨$\langle\Psi|C_{L}^{\dagger}T_{L}C_{L}|\Psi\rangle\geq 2L\langle\Psi|\Psi\rangle.$ (29)
@@
⟨$\langle\Psi|C_{L}^{\dagger}T_{L+1}C_{L}|\Psi\rangle=\langle\Psi|C_{L}^{\dagger}C_{L}|\Psi\rangle+\langle\Psi|C_{L}^{\dagger}(T_{L+1}-1)C_{L}|\Psi\rangle.$ (30)
@@
⟨$\langle\Psi|C_{L}^{\dagger}C_{L}|\Psi\rangle=2L\langle\Psi|\Psi\rangle.$ (31)
@@
The idea of the proof is shown in fig. 3. A generic product state contains domains of spins up and down of different lengths. Action of $C_{L}$ leads then to a superposition of new states which occur with multiplicities directly related to the structure of the domains of spin ups. For a state $|\Psi\rangle$ which contains a set of domains of spin up of sizes $\{m_{i}\}$ we then have
⟨$\langle\Psi|C_{L}^{\dagger}C_{L}|\Psi\rangle=\sum_{i}(m_{i}+1)^{2},$ (32)
because if by adding a spin up we enlarge a domain of length $m$ this can be achieved in $m+1$ ways and therefore contributes $(m+1)^{2}$ to the norm. For a state of zero magnetization we have a constraint ∑$\sum_{i}m_{i}=L/2$. Minimizing the right hand side under this constraint leads to $\bar{m}_{i}=1$ for $i=1,\dots,L/2$ for which ∑$\sum_{i}(\bar{m}_{i}+1)^{2}=2L$ which proves the equality (31) and in turn inequality (29).
@@
Figure 3: Illustration of contributions to ⟨$\langle\Psi|C_{L}^{\dagger}C_{L}|\Psi\rangle$. For $|\Psi\rangle$ we choose a zero magnetization state without any translational symmetry. Action of $C_{L}$ inserts spin up (in red) at $6$ different positions leading to a superposition of $3$ different states. Value of ⟨$\langle\Psi|C_{L}^{\dagger}C_{L}|\Psi\rangle$ is then directly related to the multiplicity of those states. In the example given we find ⟨$\langle\Psi|C_{L}^{\dagger}C_{L}|\Psi\rangle=2^{2}+3^{2}+1=14\geq 2L$ taking ⟨$\langle\Psi|\Psi\rangle=1$. The multiplicities are directly related to the lengths of the domains of spin ups present in the seed state. Denoting by $m$ the length of a domain to which we add a spin, the multiplicity of the resulting state is $m+1$. This is valid also for $m=0$ when we insert the spin up between the spin downs as on the right hand side of the figure.
@@
$|\Psi_{0}\rangle=|\!\uparrow\downarrow\uparrow\downarrow\cdots\rangle+|\!\downarrow\uparrow\downarrow\uparrow\cdots\rangle,$ (33)

### v3

# Nature of Spinons in 1D Spin Chains

@@ Abstract
Figure 1: Cartoon views of collective excitations in spin-$\frac{1}{2}$ Heisenberg chains. Panel (a): canonical picture arising after flipping one spin in the Heisenberg FM ground state (left subpanel) and triggering a propagating magnon (middle subpanel) which is understood as a free mobile particle in the vacuum (right subpanel). Panel (b): canonical picture arising after flipping one spin in the Heisenberg AF ground state (left subpanel) and triggering two propagating spinons (middle subpanel) which are usually approximately understood as two mobile domain walls in a Néel AF (right subpanel). Panel (c): picture introduced in this paper, arising after inserting an additional spin in the Heisenberg AF ground state (left subpanel) and triggering a propagating spinon (middle subpanel) which can be understood as a mobile spin flowing in the valence-bond solid (right subpanel). Flipped or inserted spins are shown in red; entanglement between two or more sites is shown by a blue-coloured hatched area.
@@
An important feature of the construction, further discussed below, is that $|\Psi(m)\rangle$ is not normalized. Denoting by $|\bar{\Psi}(m)\rangle$ its normalized counterpart, we find the first two moments of the local spin operators to be $\langle\bar{\Psi}(m)|\sigma_{j}^{z}|\bar{\Psi}(m)\rangle=\frac{\delta_{jm}}{2}$, $\langle\bar{\Psi}(m)|(\sigma_{m}^{z})^{2}|\bar{\Psi}(m)\rangle=\frac{1}{4}$, and for $j\neq m$: $\langle\bar{\Psi}(m)|(\sigma_{j}^{z})^{2}|\bar{\Psi}(m)\rangle=\langle\Psi_{0}|(\sigma_{1}^{z})^{2}|\Psi_{0}\rangle$, where in the last equality we used the translational invariance of the ground state. This result can be visualized as a spin up frozen at the $m^{th}$ site of the $L+1$-site chain, in the sea of fluctuating spins on all other sites. These fluctuations are identical to the ground state fluctuations of the $L$-site chain, by construction. This picture is in accordance with the concept of elementary excitations which are created from the ground state, and inherit rather than rebuild the correlations present in it.
To define a dispersion relation we introduce a Fourier transform with momentum $q=\frac{2\pi n}{L+1}$, $-L/2\leqslant n\leqslant L/2$, $n\in\mathbb{Z}$:
@@
In our procedure, the correlations present in the spinon state are inherited from the ground state. This leads to the natural question as to which correlations, of all those present in the exact ground state, are most relevant to the spinon. To answer this question we consider two approximations to the ground state: (i) the lowest energy classical configuration, i.e. the Néel state, and (ii) the simplest valence-bond solid.
@@
The corresponding $\varepsilon(q)$, $N(q)$ are shown in Fig. 2(c1-c2). The dispersion relation is qualitatively similar to the spinon dispersion (black solid line). Quantitatively, the value at $q=0$ is underestimated and the value at $q=\frac{\pi}{2}$ is negative, reflecting the fact that the valence-bond state is an approximation. Indeed, its average energy per site is $-0.3750$, significantly larger than the exact ground state energy per site of $-0.4438$ Auerbach (1994).
@@
Dispersion linear in momentum: The other notable feature of the spinon dispersion is its linear dependence on $q$ in the low-energy limit. It can be understood as a result of (i) the vanishing compact support for $|q|>\pi/2$ that is due to ground state entanglement (see above), and (ii) the $\propto\cos q$ dispersion relation arising from nearest neighbor hopping of the spinon, discussed next.
@@
Outlook: Following the publicly available draft of this paper, the procedure to excite single spinons was just tested experimentally using STM on a fabricated nanographene Heisenberg spin chain Zhao et al. (2025). In particular, the Authors of  Zhao et al. (2025) managed to observe a single spinon standing wave in an STM experiment. This not only confirms the findings of this paper but also shows that it can indeed be experimentally implemented. Finally, we suggest that also particular pump-probe experimental techniques may be able to implement this procedure, for more details see Sec. IV of SM .
The postulated procedure is universal and can be applied to any 1D spin model, cf. Sec. III of SM for a short overview. In particular, it can be used to study excitations of the 1D FM chain, see Fig. 2(a-b) of SM : in this case a single magnon is excited in the procedure. This further shows that the spinon is not exotic because it can be excited in quite a special manner but it is the entanglement nature of the antiferromagnetic ground state that is crucial here. We suggest that in the future, this procedure can be used to explore the nature of the elementary excitations in frustrated spin systems, such as the recently discussed problem of marginally irrelevant interactions between spinons in $J_{1}-J_{2}$ spin chains Keselman et al. (2020). Last but not least, our understanding is that the above-mentioned experimental techniques can in principle be carried out also in higher dimensions. This way the nature of single-spinon excitations beyond 1D can be revealed.
@@
Acknowledgements: Authors acknowledge the support of National Science Centre in Poland under the projects 2016/22/E/ST3/00560 (K.W.), 2024/55/B/ST3/03144 (K.W), 2018/31/B/ST3/03758 (T.K.), 2018/31/D/ST3/03588 (M.P.), 2022/47/B/ST2/03334 (M.P.). K.W. acknowledges support from the Excellence Initiative of the University of Warsaw (the ‘mikrogrant’ programme). M.B. acknowledges support from the Canada First Research Excellence Fund and the Natural Sciences and Engineering Research Council of Canada. We thank Federico Becca, Akira Furusaki and Karlo Penc for stimulating discussions and all the organisers of the KITP program “A New Spin on Quantum Mangets” for allowing us to present this work during the program. We thank the Referee for bringing Ref. Talstra and Strong (1997) to our attention, Karlo Penc for pointing out references  Sorella and Parola (1992); Penc et al. (1997); Penc and Serhan (1997), and Akira Furusaki for pointing out references  Matveev et al. (2007a, b). This research was supported in part by grant NSF PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP). This research was carried out with the support of the Interdisciplinary Centre for Mathematical and Computational Modelling at University of Warsaw (ICM UW) under grant no. G73-29.
@@ I Exact result for the $4$-site Heisenberg chain: relation between the norm $N(q)$ and entanglement
$|\Psi_{0}\rangle=\frac{|\!\uparrow\downarrow\rangle-|\!\downarrow\uparrow\rangle}{\sqrt{2}}\otimes\frac{|\!\uparrow\downarrow\rangle-|\!\downarrow\uparrow\rangle}{\sqrt{2}}+{\rm translation}.$ (SI.1)
@@
$|\Psi(q)\rangle=\frac{1+e^{iq}-\frac{e^{3iq}}{2}}{\sqrt{3}}|A\rangle-\frac{1+e^{iq}+e^{2iq}}{2\sqrt{3}}|B\rangle$ (SI.2)
with $|A\rangle=|\!\!\uparrow\uparrow\downarrow\uparrow\downarrow\rangle+({\rm translations})$ and $|B\rangle=|\!\!\uparrow\uparrow\uparrow\downarrow\downarrow\rangle+({\rm translations})$. All the states in $|A\rangle$ and $|B\rangle$ are orthonormal to each other. The resulting norm:
$N(q)=1+\cos q-\frac{1}{6}\cos(2q)-\frac{1}{3}\cos(3q)$ (SI.3)
@@
Having obtained an exact expression for $N(q)$, (SI.3), we are now ready to investigate the origin of its two main striking properties: (i) a relative enhancement of the $4$-site Heisenberg chain norm, with a pronounced maximum for the small momenta $q\lesssim\pi/2$, w.r.t. the norm obtained once a spin is added to the Neel antiferromagnet, that is w.r.t. $N(q)=1+\cos(q)$ (see main text); (ii) a relative suppression of the $4$-site Heisenberg chain norm, with a pronounced minimum for the small momenta $q\gtrsim\pi/2$, w.r.t. the norm obtained once a spin is added to the Neel antiferromagnet, that is w.r.t. $N(q)=1+\cos(q)$ (see main text).
To this end, let us first note that a sum rule $\sum_{0<q<\pi}N(q)=\pi$ means that once we e.g. understand the relative enhancement for some momenta, then the relative suppression for the other momenta follows. Therefore, we concentrate on investigating the former effect.
A quick inspection of Eq. (SI.3) shows that the term responsible for obtaining such a relative enhancement in $N(q)$ for small momenta (and, conversely, also a relative suppression for large momenta) is the term $\propto-\cos(3q)$. A closer look at the procedure of adding an extra spin shows that, in order to obtain such a contribution to the norm, there must exist two product states in the ground state which contribute to $|\Psi_{0}\rangle$ with opposite sign and that they must be identical to each other after adding an extra spin in our procedure, modulo a phase factor $e^{3iq}$. This is fulfilled for instance by such two contributions to the exact ground state $|\Psi_{0}\rangle$ that, upon adding an extra spin site at particular location, lead to the same states which differ by $e^{3iq}$ in our procedure:
@@
$-\frac{1}{2}|\uparrow\uparrow\downarrow\downarrow\rangle\quad$ $\rightarrow\quad-\frac{1}{2}e^{3iq}|\uparrow\uparrow\downarrow\tilde{\uparrow}\downarrow\rangle,$ (SI.5)
@@
Identifying the type of product states in $|\Psi_{0}\rangle$ that lead to the relative enhancement of the norm $N(q)$ for small momenta, allows us to intuitively understand the origin of this phenomenon: namely, it follows from the valence bond nature of the ground state—the two ‘fluctuating singlets’ allow for the onset of ferromagnetic domains with opposite magnetisation and a (relative) negative phase contribution to the ground state, see Eq. (SI.5). Note that the relative negative phase is just as crucial as the existence of ferromagnetic domains and that such a phase difference can only be realised once $|\Psi_{0}\rangle$ is not a product state. Altogether, this shows how specific properties of the norm $N(q)$ of the state constructed by adding one extra spin can be related to the entangled nature of the spin chain ground state $|\Psi_{0}\rangle$.
@@ II The norm $N(0)$ as an entanglement witness
$|\Psi(q=0)\rangle=\frac{1}{\sqrt{L+1}}\bar{C}_{L}|\Psi_{0}\rangle.$ (SII.8)
@@
$N(0)=\frac{1}{L+1}\frac{\langle\Psi_{0}|\bar{C}_{L}^{\dagger}\bar{C}_{L}|\Psi_{0}\rangle}{\langle\Psi_{0}|\Psi_{0}\rangle},$ (SII.9)
@@
$T_{L}|\Psi_{1}\rangle=T_{L}|\Psi_{2}\rangle,\qquad|\Psi_{i}\rangle\in[|\Psi_{0}\rangle].$ (SII.10)
@@
$\mathcal{N}_{\Psi_{0}}=\frac{1}{L+1}\frac{\langle\Psi|T_{L}^{\dagger}\bar{C}_{L}^{\dagger}\bar{C}_{L}T_{L}|\Psi\rangle}{\langle\Psi|T_{L}^{\dagger}T_{L}|\Psi\rangle}.$ (SII.11)
From the property (SII.10) it follows that $\mathcal{N}$ is the same for any $|\Psi\rangle$ in the equivalance class $[|\Psi_{0}\rangle]$ of $|\Psi_{0}\rangle$. This implies that to compute $\mathcal{N}$ for a translationally invariant state we can identify one of its seeds and then use equation (SII.11). This equation still explicitly symmetrizes over all translations but we will now rewrite in a form more suitable for further computations. In this process the above mentioned symmetry becomes less obvious.
@@
Using the properties of operators $C_{L}$ and $T_{L}$ we can prove a simple property of $\langle\Psi_{0}|\bar{C}_{L}^{\dagger}\bar{C}_{L}|\Psi_{0}\rangle$ that, due to the translational invariance, position of one of the insertions can be fixed,
$\langle\Psi_{0}|\bar{C}_{L}^{\dagger}\bar{C}_{L}|\Psi_{0}\rangle=\langle\Psi|T_{L}\bar{C}_{L}^{\dagger}\bar{C}_{L}T_{L}|\Psi\rangle=\langle\Psi|C_{L}^{\dagger}T_{L+1}T_{L+1}c_{m}T_{L}|\Psi\rangle=(L+1)\langle\Psi|\bar{C}_{L}^{\dagger}T_{L+1}c_{m}T_{L}|\Psi\rangle$ (SII.12)
@@
$\mathcal{N}=\frac{1}{L}\frac{\langle\Psi|T_{L}\bar{C}_{L}^{\dagger}c_{0}T_{L}|\Psi\rangle}{\langle\Psi|T_{L}|\Psi\rangle}.$ (SII.13)
@@
$|\Psi\rangle=\bigotimes_{i=1}^{k}|\psi\rangle,$ (SII.15)
@@
$\mathcal{N}_{\Psi_{0}}\geq N_{\psi_{0}}.$ (SII.16)
Assuming (SII.14) we have $T_{L}|\Psi\rangle=kT_{m}|\Psi\rangle$. We then find
@@
$\geq\frac{1}{m}\frac{\langle\Psi|T_{m}\bar{C}_{m}^{\dagger}c_{0}T_{m}|\Psi\rangle}{\langle\Psi|\Psi\rangle}=\frac{1}{m}\frac{\langle\psi|T_{m}\bar{C}_{m}^{\dagger}c_{0}T_{m}|\psi\rangle}{\langle\psi|\psi\rangle}=\mathcal{N}_{\psi_{0}},$ (SII.17)
@@
This proves that $\mathcal{N}$ for a seed $|\Psi\rangle$ which is a product state can be bounded by $\mathcal{N}$ of the elementary cell $|\psi\rangle$ of the seed. This is important because if we now want to find further bounds on $\mathcal{N}$ we can focus on states which do not have any invariance under translations which simplifies the analysis. To this end, we observe that for state $|\Psi_{0}\rangle$ such that its seed $|\Psi\rangle$ does not have any translational symmetry the formula (SII.11) can be reduced to
$\mathcal{N}_{\Psi_{0}}=\frac{1}{L}\frac{\langle\Psi|C_{L}^{\dagger}T_{L+1}C_{L}|\Psi\rangle}{\langle\Psi|\Psi\rangle},$ (SII.18)
where we used $T_{L}^{2}=LT_{L}$ and due to lack of the translational symmetry $\langle\Psi|T_{L}|\Psi\rangle=\langle\Psi|\Psi\rangle$.
@@
$\langle\Psi|C_{L}^{\dagger}T_{L}C_{L}|\Psi\rangle\geq 2L\langle\Psi|\Psi\rangle.$ (SII.19)
@@
$\langle\Psi|C_{L}^{\dagger}T_{L+1}C_{L}|\Psi\rangle=\langle\Psi|C_{L}^{\dagger}C_{L}|\Psi\rangle+\langle\Psi|C_{L}^{\dagger}(T_{L+1}-1)C_{L}|\Psi\rangle.$ (SII.20)
@@
$\langle\Psi|C_{L}^{\dagger}C_{L}|\Psi\rangle=2L\langle\Psi|\Psi\rangle.$ (SII.21)
@@
The idea of the proof is shown in fig. S1. A generic product state contains domains of spins up and down of different lengths. Action of $C_{L}$ leads then to a superposition of new states which occur with multiplicities directly related to the structure of the domains of spin ups. For a state $|\Psi\rangle$ which contains a set of domains of spin up of sizes $\{m_{i}\}$ we then have
$\langle\Psi|C_{L}^{\dagger}C_{L}|\Psi\rangle=\sum_{i}(m_{i}+1)^{2},$ (SII.22)
because if by adding a spin up we enlarge a domain of length $m$ this can be achieved in $m+1$ ways and therefore contributes $(m+1)^{2}$ to the norm. For a state of zero magnetization we have a constraint $\sum_{i}m_{i}=L/2$. Minimizing the right hand side under this constraint leads to $\bar{m}_{i}=1$ for $i=1,\dots,L/2$ for which $\sum_{i}(\bar{m}_{i}+1)^{2}=2L$ which proves the equality (SII.21) and in turn inequality (SII.19).
@@
Figure S1: Illustration of contributions to $\langle\Psi|C_{L}^{\dagger}C_{L}|\Psi\rangle$. For $|\Psi\rangle$ we choose a zero magnetization state without any translational symmetry. Action of $C_{L}$ inserts spin up (in red) at $6$ different positions leading to a superposition of $3$ different states. Value of $\langle\Psi|C_{L}^{\dagger}C_{L}|\Psi\rangle$ is then directly related to the multiplicity of those states. In the example given we find $\langle\Psi|C_{L}^{\dagger}C_{L}|\Psi\rangle=2^{2}+3^{2}+1=14\geq 2L$ taking $\langle\Psi|\Psi\rangle=1$. The multiplicities are directly related to the lengths of the domains of spin ups present in the seed state. Denoting by $m$ the length of a domain to which we add a spin, the multiplicity of the resulting state is $m+1$. This is valid also for $m=0$ when we insert the spin up between the spin downs as on the right hand side of the figure.
@@
$|\Psi_{0}\rangle=|\!\uparrow\downarrow\uparrow\downarrow\cdots\rangle+|\!\downarrow\uparrow\downarrow\uparrow\cdots\rangle,$ (SII.23)